# Resolving Build Conflicts via Example-Based and Rule-Based Program Transformations

### v1

@@ 3.4. Rule-based Transformation (BuCoR-R)
BuCoR-R resolves conflicts via rule-based program transformations. To define practical resolution rules or patterns, we manually inspected the rules and data observed by prior empirical studies (Shen et al., 2019; Da Silva et al., 2022), and expanded those rule sets to generalize rules across similar but different conflict types. Table 2 shows the resulting 16 conventional patterns we implemented in BuCoR-R; each pattern resolves one type of frequent conflicts. Among the 16 conflict types, only 6 types involve symbol renaming (C1, C6, C7, C11, C13, C15). As prior work only resolves conflicts caused by symbol renaming, they cannot handle 10 of the conflict types BuCoR-R focuses on.
@@ 4.1. Dataset
We constructed our evaluation dataset by reusing the datasets of prior work (Towqir et al., 2023; Da Silva et al., 2022; Shen et al., 2019) with our best effort. Given a merging scenario with at least one known build conflict, we decided whether it is reusable based on the following criteria:
@@ 4.2. Metrics
$C=\dfrac{\text{\# of conflicts with at least one resolution generated}}{\text{Total \# of known conflicts}}$
@@
A resolution is correct if the resolved version is semantically equivalent to the ground truth—developers’ resolution recorded in $m$.
@@ 4.3. RQ1: BuCoR Resolution Coverage
/* Left Version: PathItem.java */
@@
/* Right Version: Reader.java */
@@
/* Left Version: JedisCluster.java */
@@
/* Right Version: JedisCluster.java */
@@
/* Left Version: */
@@
/* Right Version: ClientMultiMapProxy.java */
@@
/* Left Version: RedisClient.java */
@@
/* Right Version: RedisClient.java */
@@
BuCoR-R is applicable to many cases where (1) BuCoR-E does not work and (2) the needed resolutions are simple or conventional, such as the case in Listing 2. However, for unconventional conflicts (see C17–C23 in Table 3) that involve entity deletion, parameter addition/deletion, or type change, BuCoR-R generates nothing because there is no typical, generally accepted way to handle them. For example, Listing 3 shows a conflict from Jedis (jed, 2025). This conflict is of C20–Field: remove def vs. add use, where $l$ removes field timeout and $r$ introduces a new reference to timeout. Such a conflict may get resolved by adding back the field def, replacing the added use with something else, or simply removing that use; nevertheless, there is no standard solution or typical resolution pattern followed by developers or documented in literature. Therefore, BuCoR-R did not handle it, while BuCoR-E resolved it in a project-specific way by referring to edit examples. It removed field use as below:
@@ 4.4. RQ2: BuCoR Resolution Accuracy
As shown in Table 4, 21 of the 28 resolutions produced by BuCoR-E are correct, leading to 75% accuracy; 29 of the 51 resolutions output by BuCoR-R are correct, achieving 57% accuracy. Overall, BuCoR obtained 63% accuracy, having 50 of the 79 generated resolutions to be correct. These measurements also show that the two strategies complement each other. Among the 14 cases where both strategies suggest resolutions, BuCoR-E and BuCoR-R correctly resolved 13 and 7 cases, respectively; the correct resolutions overlap in 7 cases. Thus, BuCoR resolved 43 cases correctly.
@@
/* Left Version: TypeUtils.java */
@@
/* Right Version: TypeUtils.java */
@@
BuCoR-R sometimes suggested incorrect resolutions, because the frequently applied resolution patterns within it do not always match developers’ practices. For instance, the conflict in Listing 6 has $l$ remove an import and $r$ add a reference to the removed imported class. BuCoR-R resolved this conflict by adding back the removed import. However, developers implemented a replacement method, to avoid the removed dependency. Additionally, for conflicts due to class-hierarchy changes, BuCoR-R modified subclasses to match changes in super classes, while developers sometimes adapted super classes to subclasses.
The 43 conflicts correctly resolved by BuCoR fall into 15 categories, implying that BuCoR can properly resolve diverse conflicts. There are four categories of conflicts for which BuCoR produced some resolutions but none of them are correct: C2, C9, C12, and C18. Three of the categories (C2, C9, and C12) are about mismatches between super types and sub types; two of the categories are caused by parameter list changes (C9 and C18). Among the seven incorrect resolutions suggested by BuCoR-E, two are partially incorrect, by presenting subsets of the needed changes; five are totally irrelevant. Among the 22 incorrect resolutions output by BuCoR-R, 6 are partially incorrect, while 16 are irrelevant.
Finding 2 (Answer to RQ2): BuCoR generated in total 50 correct resolution(s), achieving 63% (50/79) accuracy. BuCoR-E and BuCoR-R separately contributed 21 and 29 correct resolutions.
@@
First, BuCoR-R has higher coverage (58% vs. 32%) by resolving more conflicts; BuCoR-E achieves higher accuracy by having more of its resolutions match developers’ intents (75% vs. 57%). It implies that if developers (1) want to automatically resolve as many conflicts as possible in conventional ways, and (2) have good testing to detect wrong resolutions, they can rely more on BuCoR-R. If developers (1) want to automatically resolve conflicts in unconventional ways but (2) have poor support to detect wrong resolution, they can rely more on BuCoR-E. For instance, when both strategies output resolutions (see Table 1), it is likely that BuCoR-E’s result is better.
Second, the two strategies are good at resolving different types of conflicts. Among the 43 conflicts correctly handled by BuCoR, 7 conflicts are correctly resolved by both strategies. In addition to that, BuCoR-E correctly resolved 14 cases; for 8 of these cases, BuCoR-R could not suggest anything. The eight cases are related to entity removal (i.e., C17, C20, C23) and parameter-list changes (i.e., C21). It means that BuCoR-E is better at resolving conflicts in unconventional or complex ways. Meanwhile, BuCoR-R correctly resolved 22 cases, for which BuCoR-E suggested nothing. These 22 conflicts are mainly related to super-sub mismatches (C3–C4, C8, C10–C11) and duplicated entities (C14, C16). It means that BuCoR-R is better at handling simple conflicts in conventional ways.
@@ 5. Threats to Validity
Threats to External Validity. Our evaluation dataset consists of 88 real-world build conflicts, spanning 21 distinct conflict types. However, the program contexts and conflict patterns covered by this dataset may not fully capture the diversity of build conflicts in real world. BuCoR-R includes 16 well-defined rules of handling frequently occurring conflicts. Although they cover all the typical conflict-resolution strategies we are aware of, they may not include all the frequently applied strategies in reality. In the future, to enhance our research representativeness, we will (1) expand our dataset to include more merging scenarios, and (2) enlarge the ruleset of BuCoR-R whenever possible.
@@ 6.1. Automated Software Merge
Detectors of Build Conflicts (Wuensche et al., 2020b; Towqir et al., 2023). The existing tools both statically analyze branch versions, reason about branch edits, and contrast extracted edits with predefined patterns to detect conflicts. However, neither tool automatically resolves build conflicts.
@@
Resolvers of Build Conflicts (Sung et al., 2020; Zhang et al., 2022). MrgBldBrkFixer (Sung et al., 2020) compares the ASTs of C++ files, to resolve conflicts related to renaming changes. Gmerge (Zhang et al., 2022) applies few-shot learning to GPT-3, to resolve conflicts due to renaming. They are analogous to BuCoR-E, because they all infer and apply patterns by analyzing exemplar edits. We did not empirically compare BuCoR with either tool, as they are close-sourced tools to handle conflicts in C/C++ code.
@@ 6.2. Empirical Studies of Merge Conflicts
Relationship between Merge Conflicts and Software Maintenance (Estler et al., 2014; Ahmed et al., 2017; Mahmoudi et al., 2019). Estler et al. (Estler et al., 2014) surveyed 105 student developers, to study the relationship between awareness (i.e., knowing “who’s changing what”) and merge conflicts. Ahmed et al. (Ahmed et al., 2017) explored the effect of bad design (code smells) on merge conflicts. Mahmoudi et al. (Mahmoudi et al., 2019) studied the relation between merge conflicts and 15 popular refactoring types. These researchers showed that merge conflicts are widespread, although they studied textual conflicts.

### v2

@@ 3.4. Rule-based Transformation (BuCoR-R)
BuCoR-R resolves conflicts via rule-based program transformations. To define practical resolution rules or patterns, we manually inspected the rules and data observed by prior empirical studies (Sung et al., 2020; Da Silva et al., 2022), and expanded those rule sets to generalize rules across similar but different conflict types. Table 2 shows the resulting 16 conventional rules we implemented in BuCoR-R; each rule resolves one type of frequent conflicts. Among the 16 rules, only 6 rules involve symbol renaming (C1, C6, C7, C11, C13, C15). As prior work only resolves conflicts caused by symbol renaming, they cannot handle 10 of the conflict types BuCoR-R focuses on. Only 8 of the 16 rules are clearly described by prior work: C1, C3, C5–C6, C8, C14–C16. We derived the remaining rules by generalizing across related conflict types.
@@ 4.1. Dataset
We constructed our evaluation dataset by reusing the datasets of prior work (Towqir et al., 2023; Da Silva et al., 2022) with our best effort. Given a merging scenario with at least one known build conflict, we decided whether it is reusable based on the following criteria:
@@ 4.2. Metrics
[…]
@@
A resolution is correct if the resolved version is semantically equivalent to the ground truth—developers’ resolution recorded in $m$. Two authors independently inspected BuCoR’s resolution to examine semantic alignment, and had a discussion whenever they disagree.
@@ 4.3. RQ1: BuCoR Resolution Coverage
/* Left Version: PathItem.java adds a method call to setRef(…) */
@@
/* Right Version: Reader.java renames setRef(…) */
@@
/* Left Version: JedisCluster.java removes a field */
@@
/* Right Version: JedisCluster.java adds a reference to field timeout */
@@
/* Left Version moves a method between classes */
@@
/* Right Version: ClientMultiMapProxy.java adds calls to the original method */
@@
/* Left Version: RedisClient.java adds a call to one of its constructors */
@@
/* Right Version: RedisClient.java adds a parameter to one of its constructors */
@@
BuCoR-R is applicable to many cases where (1) BuCoR-E does not work and (2) the needed resolutions are conventional, such as the case in Listing 2. However, for conflicts which have no conventional resolution method recommended by literature (see C17–C23 in Table 3), BuCoR-R generates nothing because there is no typical, generally accepted way to handle them. For example, Listing 3 shows a conflict from Jedis (jed, 2025). This conflict is of C20–Field: remove def vs. add use, where $l$ removes field timeout and $r$ introduces a new reference to timeout. Such a conflict may get resolved by adding back the field def, replacing the added use with something else, or simply removing that use; nevertheless, there is no standard solution or typical resolution pattern followed by developers or documented in literature. Therefore, BuCoR-R did not handle it, while BuCoR-E resolved it in a project-specific way by referring to edit examples. It removed field use as below:
@@ 4.4. RQ2: BuCoR Resolution Accuracy
As shown in Table 4, 21 of the 28 resolutions produced by BuCoR-E are correct, leading to 75% accuracy; 20 of the 51 resolutions output by BuCoR-R are correct, achieving 39% accuracy. Overall, BuCoR obtained 52% accuracy, having 41 of the 79 generated resolutions to be correct. These measurements also show the two strategies to complement each other. Among the 14 cases where both strategies suggest resolutions, BuCoR-E and BuCoR-R correctly resolved 13 and 7 cases, respectively; the correct resolutions overlap in 7 cases. Thus, BuCoR resolved 34 cases correctly.
@@
/* Left Version: TypeUtils.java removes an import */
@@
/* Right Version: TypeUtils.java adds a reference to the imported class */
@@
BuCoR-R suggested incorrect resolutions for three reasons. First, its resolution rules do not always match developers’ practices. For instance, the conflict in Listing 6 has $l$ remove an import and $r$ add a reference to the removed imported class. BuCoR-R resolved this conflict by adding back the removed import. However, developers implemented a replacement method, to avoid the removed dependency. Second, for conflicts due to class-hierarchy changes, BuCoR-R modified subclasses to match changes in super classes, while developers sometimes adapted super classes to subclasses. Third, when conflict resolution requires implementation of a new method or updates of a method’s parameter/return type, BuCoR-R can only change the method signature, without customizing method implementation accordingly.
The 34 conflicts correctly resolved by BuCoR fall into 12 categories, implying that BuCoR can properly resolve diverse conflicts. There are seven conflict categories for which BuCoR produced some resolutions but none of them are correct: C2–C4, C8–C9, C12, and C18. Six of the categories (C2–C4, C8–C9, and C12) are about mismatches between super types and sub types; three of the categories are caused by parameter list changes (C3, C9 and C18). Among the seven incorrect resolutions output by BuCoR-E, two are partially incorrect, by presenting subsets of the needed changes; five are totally irrelevant. Among the 31 incorrect resolutions output by BuCoR-R, 15 are partially incorrect, while 16 are irrelevant.
Finding 2 (Answer to RQ2): BuCoR generated in total 41 correct resolution(s), achieving 52% (41/79) accuracy. BuCoR-E and BuCoR-R separately contributed 21 and 20 correct resolutions.
@@
First, BuCoR-R has higher coverage (58% vs. 32%) by resolving more conflicts; BuCoR-E achieves higher accuracy by having more of its resolutions match developers’ intents (75% vs. 39%). It implies that if developers (1) want to automatically resolve as many conflicts as possible in conventional ways, and (2) have good testing to detect wrong resolutions, they can rely more on BuCoR-R. If developers (1) want to automatically resolve conflicts in unconventional ways but (2) have poor support to detect wrong resolution, they can rely more on BuCoR-E. For instance, when both strategies output resolutions (see Table 1), it is likely that BuCoR-E’s result is better.
Second, the two strategies are good at resolving different types of conflicts. Among the 34 conflicts correctly handled by BuCoR, 7 conflicts are correctly resolved by both strategies. In addition to that, BuCoR-E correctly resolved 14 cases; for 8 of these cases, BuCoR-R could not suggest anything. The eight cases are related to entity removal (i.e., C17, C20, C23) and parameter-list changes (i.e., C21). It means that BuCoR-E is better at resolving conflicts in unconventional ways. Meanwhile, BuCoR-R correctly resolved 13 cases, for which BuCoR-E suggested nothing. These 13 conflicts are mainly related to super-sub mismatches (C10–C11) and duplicated entities (C14, C16). It means that BuCoR-R is better at handling conflicts in conventional ways.
@@ 5. Threats to Validity
Threats to External Validity. Our evaluation dataset consists of 88 real-world build conflicts, spanning 21 distinct conflict types. However, the program contexts and conflict patterns covered by this dataset may not fully capture the diversity of build conflicts in real world. BuCoR-R includes 16 well-defined rules of handling frequently occurring conflicts. Although they cover all the typical conflict-resolution strategies mentioned by prior work (Da Silva et al., 2022; Shen et al., 2022), they may not include all the frequently applied strategies in reality. In the future, to enhance our research representativeness, we will (1) expand our dataset to include more merging scenarios, and (2) enlarge the ruleset of BuCoR-R whenever possible.
@@ 6.1. Automated Software Merge
Detectors of Build Conflicts (Wuensche et al., 2020; Towqir et al., 2023). The existing tools both statically analyze branch versions, reason about branch edits, and contrast extracted edits with predefined patterns to detect conflicts. However, neither tool automatically resolves build conflicts.
@@
Resolvers of Build Conflicts (Sung et al., 2020; Zhang et al., 2022). MrgBldBrkFixer (Sung et al., 2020) compares the ASTs of C++ files, to resolve conflicts related to renaming changes. Gmerge (Zhang et al., 2022) applies few-shot learning to GPT-3, to resolve conflicts due to renaming. They are analogous to BuCoR-E, because they all infer and apply patterns by analyzing exemplar edits. We did not empirically compare BuCoR with either tool, as they are closed-source and target C/C++ conflicts.
To assess the effectiveness of large language models (LLMs) on our dataset, we recently conducted an extra experiment. We prompted OpenAI’s o4-mini with the full diffs of (base, left) and (base, right) of each conflicting merging scenario. Two authors manually validated o4-mini’s outputs. Due to token limit, o4-mini was applicable to only 48 scenarios. It detected 117 conflicts, of which 63 were correct. It correctly resolved 25 of the conflicts—substantially worse than BuCoR.
@@ 6.2. Empirical Studies of Merge Conflicts
Relationship between Merge Conflicts and Software Maintenance (Estler et al., 2014; Ahmed et al., 2017; Mahmoudi et al., 2019; Vale et al., 2023). Estler et al. (Estler et al., 2014) surveyed 105 student developers, to study the relationship between awareness (i.e., knowing “who’s changing what”) and merge conflicts. Ahmed et al. (Ahmed et al., 2017) explored the effect of bad design (code smells) on merge conflicts. Mahmoudi et al. (Mahmoudi et al., 2019) studied the relation between merge conflicts and 15 popular refactoring types. These researchers showed that conflicts are widespread, although they studied textual conflicts.
@@ 7. Conclusion
Acknowledgements.
We thank all reviewers for their valuable feedback. This work was funded by NSF-1845446.